OpenStudy (anonymous):

What does it mean for a number to be the "solution" to an equation? Imagine you are teaching a fellow student how to solve: 4 − (x + 2) = 6 − 5(x + 4) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

for a number to be the soluton to an equation when it is plugged into the equation it must solve it

OpenStudy (anonymous):

i dnt exactly know how your teacher may want this done but u should simply the eqn then solve for x...so u wld have 4-x-2=6-5x-20 next line would be rearange 4-2-x=6-20-5x next line wld be 2-x=-14-5x next line wld be 2+14=-5x+x next line 16=-4x divide bothe sides by -4 and x= -4

OpenStudy (anonymous):

4-((-4)+2) should equal 6-5((-4)+4) 4-(-2) 6-5 (0) .....5x0=0 so 6-0 6 6

OpenStudy (precal):

 set them up as 2 equations to graph and the solution will be the cross point